Use g = 980.35 cm/s^2 foe this problem. a) A 68.44
Answer #1

Data Given

g = 980.35 m/s2, Horizontal speed u = 601.78 cm/s Vertical height h = 104.53 cm

Part A) Lets first calculate time of flight in case of horizontal projectile

T = \sqrt{\frac{2h}{g} } =\sqrt{\frac{2\times 104.53}{980.35} }= 0.46 s

Now

Part B) Initial momentum of the ball will be

p_{i,ball} = m.u = 68.44g\times 601.78 m/s = 41185.82 g.cm/s

Part C) Velocity of the ball before collision will be

v = \sqrt{u^{2}+(gt)^{2}}= \sqrt{(601.78)^{2}+(980.35\times 0.46 )^{2}}

v = 752 cm/s

Now using law of conservation of momentum as

mv+M.0 = (m+M)v_{f}\Rightarrow v_{f}= \frac{mv}{m+M}

v_{f}= \frac{68.44\times 752}{68.44+279.44} = 147.94 cm/s

Part D) Using conservation of energy after collision

\frac{1}{2}(m+M)v_{f}^{2}= (m+M)gh

h = \frac{v_{f}^{2}}{2g} = \frac{147.94^{2}}{2\times 980.35} = 11.16 cm

Part E)

KE before collision is

KE_{before}= \frac{1}{2}mv^{2}= \frac{1}{2}\times 68.44\times 752^{2}=1.935\times 10^{8}erg

KE_{after}= \frac{1}{2}(m+M)v_{f}^{2}= \frac{1}{2}\times (68.44+279.44)\times 147.94^{2}=3.81\times 10^{6}erg